Draw the function f(x) = pi/L x defined in the interval -L < x < L and obtain its Fourier series expansion


1
Expert's answer
2021-09-10T06:11:44-0400


Fourier series for any function f(x) defined in the interval -L < x < L is

"f(x)=\frac{a_0}{2}+\sum_{n=1}^{\infty}\left[a_ncos\left(\frac{\pi nx}{L}\right)+b_nsin\left(\frac{\pi nx}{L}\right)\right]"

where   "a_n=\frac{1}{L}\int_{-L}^{L}{f(x)cos\left(\frac{\pi nx}{L}\right)dx}" , "b_n=\frac{1}{L}\int_{-L}^{L}{f(x)sin\left(\frac{\pi nx}{L}\right)dx}"

For given function f(x) = πx/L

"a_n=\frac{\pi}{L^2}\int_{-L}^{L}xcos\left(\frac{\pi nx}{L}\right)dx=\frac{1}{\pi}\int_{-\pi}^{\pi}ycos\left(ny\right)dy=0"

"a_n=\frac{\pi}{L^2}\int_{-L}^{L}xsin\left(\frac{\pi nx}{L}\right)dx=\frac{1}{\pi}\int_{-\pi}^{\pi}ysin\left(ny\right)dy=-\frac{2}{n}cos(\pi n)+\frac{2}{\pi n^2}sin(\pi n)=\frac{2{(-1)}^{n+1}}{n}"

So the Fourier series expansion for given function is

"\frac{\pi x}{L}=2\sum_{n=1}^{\infty}{\frac{{(-1)}^{n+1}}{n}sin\left(\frac{\pi nx}{L}\right)}"